# Revisit, Extend, and Enhance Hessian-Free Influence Functions

### v2

@@ C.1 IP for measuring fairness and robustness
Similar to Figure 2, we conduct the experiments based on Eqs. (3) and (4) to demonstrate the effectiveness of our IP in measuring the sample influence on fairness and robustness. Figure 8 shows the relationship between IP and the traditional influence function, indicate that IP is a good surrogate of the traditional influence function in measuring model’s fairness and robustness as well.
C.2 Performance of IP Ensemble with different rates of removed samples
We also explore different rates of removed samples on CIFAR-10N-a. Table 4 shows the performance of IP Ensemble with different rates of removed samples. Except for not removing, there is no significant difference in remaining results.
@@ C.2 Performance of IP Ensemble with different rates of removed samples
C.3 Parameter Analysis on Dropout Rate and Ensemble Size
@@
C.4 Performance on ViT and MLP-Mixer
@@
C.5 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
Table 7 shows the time complexity of various influence-based methods. Except for the vanilla one, all other methods have linear time complexity in terms of the sample size. However, they have large divergence in real execution time.
@@ C.5 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
Appendix D Broader Impact and Limitations
Our work aims to address issues that currently hinder the applicability of influence estimation in deep learning models. By enabling influence estimation for deep models, practitioners can assess whether training samples are beneficial or detrimental to performance. As we demonstrate through extensive experiments across multiple problem settings, our proposed outlier gradient analysis approach outperforms existing baselines and can augment model performance by trimming detrimental samples. As a result, our work paves the way for significant positive societal impact, especially with the increased adoption of larger and deeper neural networks such as LLMs.
However, despite our research demonstrating superior performance and potential benefits, there are several limitations to consider. First, our method assumes a level of model and data homogeneity that might not hold in highly heterogeneous datasets or more complex model architectures. Furthermore, there is the consideration of robustness and fairness across different domains and types of data; our current evaluations, while extensive, may not cover all possible scenarios or edge cases, potentially limiting the generalizability of our findings. Finally, ethical considerations around data trimming and sample selection must be carefully managed to avoid unintended biases or negative impacts on model performance for underrepresented groups. As influence estimation becomes more integrated into model training and evaluation pipelines, ongoing research will be essential to ensure that these techniques are applied responsibly and equitably.
Due to the extensive computational demands of our experiments (see details in Appendix E), which require multiple retraining cycles to evaluate model stability and robustness, it becomes impractical for us to conduct comparative experiments on very large datasets, e.g., ImageNet. The limited computational resources at our disposal make it challenging to perform such large-scale retraining efforts, especially given the significant training time and GPU requirements. As a result, we focus our evaluation on middle-size yet representative benchmarks that allow for more feasible experimentation while maintaining the rigor of our comparative analyses.
Appendix E Code and Reproducibility
Experiments were run on Ubuntu 20.04.6 with NVIDIA RTX 4090 GPUs (24GB VRAM), CUDA 12.3, and driver 545.23.08. The code will be released upon request.

### v4

@@ A.1 IP for Measuring Fairness and Robustness
Similar to Figure 1, we conduct the experiments based on Eqs. (3) and (4) to demonstrate the effectiveness of our IP in measuring the sample influence on fairness and robustness. Figure 5 shows the relationship between IP and the traditional influence function, indicating IP is a good surrogate of the traditional influence function in measuring fairness and robustness.
@@
A.2 Performance of IP Ensemble with different rates of removed samples
In our paper, we explore different rates of removed samples on CIFAR-10N-a. Table 3 shows the performance of IP Ensemble with different rates of removed samples. Except for not removing, there is no significant difference in the remaining results.
A.3 Parameter Analysis on Dropout Rate and Ensemble Size
@@
A.4 Performance on ViT and MLP-mixer
A.5 Time Complexity and Execution Time of Various Influence Function-based Methods on Vision Datasets
@@
Table 6 shows the time complexity of various influence function-based methods. Except for the vanilla calculation, all other methods have linear time complexity in terms of the sample size. However, they have large divergence in real execution time.